OpenStudy (anonymous):

integration of 1/y^2

zepdrix (zepdrix):

\[\large \int\limits \frac{1}{y^2}\;dy \qquad =\qquad \int\limits y^{-2}\;dy\] From here just apply the `Power Rule for Integration`: ~Increase the power by 1, ~Divide by this new exponent. And of course, don't forget a constant of integration.